OpenStudy (randomecorner):

what value of p solves the eqatian? p^-5=1/243

OpenStudy (randomecorner):

i need help big time

Nnesha (nnesha):

this is like 5th time ii'm looking at this (exact same ) question

Nnesha (nnesha):

familiar with the exponent rule ? \[\huge\rm x^{-m} =\frac{ 1 }{ x^m}\]

OpenStudy (randomecorner):

nope

OpenStudy (randomecorner):

its 3 right?

Nnesha (nnesha):

x^{-m} is same as \[\huge\rm \frac{ x^{-m} }{ 1 }= \frac{ 1 }{ x^m }\] flip the fraction to convert negative to positive exponent

OpenStudy (randomecorner):

so it is 3

Nnesha (nnesha):

how did you get that ? plz explain

OpenStudy (randomecorner):

3*3*3*3*3=243

Nnesha (nnesha):

alright nice what about \[\huge\rm p^{-5}\]how would you get rid of the negative exponent ?

OpenStudy (randomecorner):

1/3

Nnesha (nnesha):

no just p^{-5}

Nnesha (nnesha):

remember flip the fraction and change the sign of exponent ?

OpenStudy (randomecorner):

wow thx

OpenStudy (randomecorner):

i feel dumb lol

Nnesha (nnesha):

well you're not

OpenStudy (randomecorner):

some of these math problems get fustrating

OpenStudy (randomecorner):

so the answer is 3 or 1/3?

Nnesha (nnesha):

well need to pay attention i'll give u an example \[\huge\rm a^{-2} = \frac{ 1 }{a^2 }\]

Nnesha (nnesha):

so \[\rm p^{-5}= ?\]

OpenStudy (randomecorner):

i was making sure because on the thing its a box to type the answer in

OpenStudy (randomecorner):

p^-5=3

Nnesha (nnesha):

or you can substitute p for 1/3 and then for 3 to check ur answer you should get equal sides at the end

Nnesha (nnesha):

how would you convert p^{-5} to positive exponent ?

OpenStudy (randomecorner):

flip the eqation

Nnesha (nnesha):

yes right how would you do that p^{-5}/1 = ?

OpenStudy (randomecorner):

-5\p=1\3

OpenStudy (randomecorner):

//

OpenStudy (randomecorner):

s in the box it is 1/3? or 3 k12 is usally not settled about that

Nnesha (nnesha):

how did you get 1/3

Nnesha (nnesha):

3* 3*3*3*3 = 243 right agree ?

OpenStudy (randomecorner):

imso confussed again sorry

OpenStudy (randomecorner):

and also 1/3 goes into 1/243 5 times

Nnesha (nnesha):

when you multiply same bases we should add their exponents so \[\huge\rm 3^1 \times 3^1 \times 3^1 \times 3^1 \times 3^1 =3^{1+1+1+1+1}\]

OpenStudy (randomecorner):

yes

Nnesha (nnesha):

so \[\huge\rm p^{-5} = \frac{ 1 }{ 3^5 }\] now move 3^5 at the numerator flip the fraction when you do that sign of the `exponent` would change

Nnesha (nnesha):

or in other words what is the reciprocal of 1/3^5 ??

OpenStudy (randomecorner):

1/15?

Nnesha (nnesha):

it's \[\huge\rm \frac{ 1 }{ 3^5 }\] 3 to the 5th power not 3 times 5

OpenStudy (randomecorner):

sorry that's what i should of wrote insted

Nnesha (nnesha):

okay so when you flip 1/3^5 what would you get ?

OpenStudy (randomecorner):

1/243

OpenStudy (randomecorner):

the answer is 3

Nnesha (nnesha):

well we changed 243 to 3^5 you should remember these exponents rule \[\huge\rm x^{-m} =\frac{ 1 }{ x^m } ~~and~~ \frac{ 1}{ 2^{-m} }= 2^m\] you have to flip the fraction to change negative to postiive exponent

Nnesha (nnesha):

i've few mints have to go

Nnesha (nnesha):

well yes \[3^{-5} = \frac{ 1 }{ 3^5 }\]

Nnesha (nnesha):

which is equal to right side